# Never-Ending Behavior-Cloning Agent for Robotic Manipulation

### v2

@@ 4.2 Comparison Performance
We present comparison results between NBAgent and other methods on Kitchen and Living Room datasets in Tabs. 1, 3, 4 and Fig. 2. As shown in Tab. 1, NBAgent significantly outperforms compared methods by $2.5\%\sim 41.4\%$ in terms of Avg metric on four NBRL settings. This indicates that NBAgent effectively addressing old skill forgetting through skill-specific and skill-shared knowledge learning. Furthermore, we presented the skill-wise comparison results in Tabs. 3 and 4. Our model exhibits the highest success rate across all manipulation skills in terms of All. metric, improving by $4.4\%\sim 36.3\%$ and $6.2\%\sim 25.6\%$ on two datasets respectively, compared with Fine-Tuning and ER. This further demonstrates the effectiveness of our model in addressing the NBRL problem. Additionally, our NBAgent is lower than ER by $6.7\%$ in terms of For metrics on Living Room 6-2 setting, which suggests the robust and significant performance of NBAgent over the whole NBRL process. Surprisingly, as shown in Tabs. 3 and 4, our NBAgent performs better than joint training methods, specifically PerAct and GNFactor, providing additional evidence to support the efficacy of our model. The additional comparison results under other settings can be found in Appendix A.3.
@@
Table 2: Comparison results in terms of success rate (%) on Living Room dataset when setting the various size of memory buff $\mathcal{M}$.
@@
4.3 Ablation Studies
Ablation on skill-shared attribute. Compared to Ours, the scores of Ours-w/oSRD on old skills are dropped by $3.8\%\sim 12.5\%$ (i.e., 1-5 and 1-6 in Tab. 1). This indicates that SRD can effectively learn skill-shared knowledge to tackle semantic overlooking on old skills, To evaluate the effectiveness of SSR module, we visualize the rendering results in Fig. 3. It suggests that SSR module can efficiently transfer skill-shared semantics of 3D scenes, thereby achieving an improvement about $0.5\%\sim 3.1\%$ in terms of Avg compared Ours-w/oSEP&SRD with ER in Tab. 1. The above mentioned demonstrates that NBAgent can significantly tackle old skill forgetting by preserving skill-shard semantics of 3D scenes and voxel representation.
Ablation on skill-specific attribute. To evaluate effectiveness of our NBAgent on learning skill-specific knowledge, we eliminate our proposed SEP module and present results in Tabs. 1, 3 and 4. Ours-w/oSRD outperforms Ours-w/oSEP&SRD on novel skills by $1.3\%\sim 11.4\%$, (i.e., 6-10 and 7-12 in Tab. 1), which demonstrates that SEP benefits our model in learning novel skills by performing skill-specific knowledge learning.
Ablation on memory size.We also explore the impact of various sizes of memory buffer $\mathcal{M}$. As shown in Tab. 2, with $|\mathcal{M}|=6$, the forgetting rate of our model notably reduced by $1.1\%\sim 10.8\%$ in comparison to $|\mathcal{M}|=4$ and $2$, respectively. This indicates that increasing the memory size significantly addresses catastrophic forgetting but also incurs a larger memory load.
Table 3: Comparison results on Kitchen under the setting of 5-1.
Skill ID 1 2 3 4 5 6 7 8 9 10 All Imp.
PerAct 96.0 77.3 53.3 30.7 37.3 52.0 2.7 18.7 4.0 73.3 44.5 $\Uparrow$ 1.5
GNFactor 92.0 60.0 70.7 12.0 46.7 52.0 5.3 22.7 10.7 70.7 44.3 $\Uparrow$ 1.7
Fine-Tuning 17.3 0.0 0.0 0.0 0.0 8.0 0.0 0.0 9.3 64.0 9.7 $\Uparrow$ 36.3
ER 90.7 54.7 37.3 54.7 28.0 60.0 5.3 0.0 24.0 62.7 41.6 $\Uparrow$ 4.4
Ours-w/oSEP&SRD 86.7 49.3 34.7 65.3 48.0 66.7 9.3 24.0 5.3 46.7 43.6 $\Uparrow$ 2.4
Ours-w/oSRD 68.0 65.3 50.7 25.3 33.3 76.0 18.7 24.0 22.7 64.0 44.8 $\Uparrow$ 1.2
[…]
@@ 4.3 Ablation Studies
Table 4: Comparison results on Living Room under the setting of 6-3.
Skill ID 1 2 3 4 5 6 7 8 9 10 11 12 All Imp.
PerAct 5.3 66.7 0.0 84.0 38.7 13.3 16.0 2.7 41.3 0.0 97.3 6.7 31.0 $\Uparrow$ 12.6
GNFactor 2.7 92.0 1.3 84.0 80.0 28.0 8.0 1.3 46.7 32.0 100 0.0 40.0 $\Uparrow$ 3.6
Fine-Tuning 8.0 18.7 0.0 0.0 0.0 0.0 0.0 0.0 0.0 68.0 100 10.7 17.1 $\Uparrow$ 25.6
ER 10.7 58.7 24.0 81.3 76.0 12.0 5.3 4.0 78.7 8.0 89.3 1.3 37.4 $\Uparrow$ 6.2
Ours-w/oSEP&SRD 13.3 74.7 28.0 76.0 52.0 8.0 9.3 12.0 53.3 25.3 96.0 2.7 37.6 $\Uparrow$ 6.0
Ours-w/oSRD 8.0 78.7 6.7 81.3 60.0 12.0 9.3 20.0 52.0 9.3 90.7 25.3 37.8 $\Uparrow$ 5.8
[…]
5 Conclusion
@@
In this paper, we explore a pioneering Never-ending Behavior-cloning Robot Learning (NBRL) problem and propose a novel NBAgent to continually learn skill-wise knowledge. Specifically, we design a skill-shared semantic rendering module and a skill-shared representation distillation module to tackle 3D scene representation overlooking on past encountered skills. Meanwhile, we propose a skill-specific evolving planner to decouple the skill-wise knowledge to effectively learning novel skills and tackling catastrophic forgetting. We develop two NBRL benchmarks to demonstrate the manipulation accuracy, which verifies the effectiveness of our NBAgent against baselines.
Appendix A Supplemental Material
This supplemental material introduces additional details not mentioned in the main paper. Overall, it mainly includes the following aspects:
(a) More Dataset details of Kitchen and Living Room in Sec. A.1.
(b) More implementation details of our proposed NBAgent in Sec. A.1.
(c) More comparison experiments in Sec. A.3.
(d) Analysis of limitations Sec. A.4.
(e) Analysis of societal impact in Sec. A.5.
Table 5: Some important notations and definitions used in this paper.
Notations Definitions Notations Definitions
$\mathcal{T}$ The skill data stream $\mathbf{Y}_{c}$ The input image with ground truth color
$\mathcal{T}^{m}$ The $m$-th continual skill learning task $\mathbf{Y}_{c,t}$ A noisy image
$\mathcal{D}^{m}_{i}$ The demonstration of $i$-th skill in $m$-th task $\mathbf{M}$ The 2D semantic map
$\mathbf{a}$ Action space $\Theta_{u}$ The Stable Diffusion model
$\mathbf{r}$ Structured observation $\mathcal{E}_{v}$ The VAE encoder for image $\mathbf{Y}_{c}$
$\mathbf{l}$ Language instruction $\mathbf{F}_{s}$ The rendered vision-language semantic feature
$\mathbf{Q}^{m}$ The Q-function head in $m$-th continual skill learning task $\hat{\mathbf{F}_{s}}$ The diffusion vision-language semantic feature
$\Theta^{m}_{p}$ The perceiver model in $m$-th continual skill learning task $t$ The diffusion process step
$\Theta^{m}_{n}$ The NeRF model in $m$-th continual skill learning task $\mathbf{l}_{p}$ The language prompt modified from $\mathbf{l}$
$\mathcal{F}_{\Theta^{m}_{n}}$ The neural rending function of $\Theta^{m}_{n}$ $\mathcal{E}_{c}$ The language encoder from CLIP
$\mathbf{v}$ The 3D voxel observation $\mathbf{l}_{s}$ A skill-specific language semantic embedding
$\mathbf{v}_{s}$ The 3D voxel feature $\mathcal{B}$ The adaptive language semantic bank
$\mathbf{x}$ The 3D input point $\mathcal{C}$ The cosine similarity matrix
$\sigma$ Differential density $\mathrm{nonzero}(\cdot)$ An operation employed to calculate the number of nonzero vectors in $\mathcal{B}$
$\mathbf{d}$ The unit viewing direction $h$ A skill-wise code
$\mathbf{r}$ Camera ray $\mathbf{S}$ Dynamic skill-specific latent space
$\mathcal{R}$ The set of all camera rays $\mathbf{F}_{c}$ The cross-attention feature
$\mathbf{o}$ Camera origin $\rho$ The softmax function
$\mathbf{c}$ The color of RGB $\mathbf{e}$ The encoded embeddings of input language instruction
$\mathbf{s}$ The semantic feature $\mathbf{W}$ The MLP projection layer in Transformer block
$\mathbf{C}$ A RGB image $\mathbf{k}^{m}_{j}$ The $j$-th keyframe in a skill, is same with $\mathbf{k}^{m,i}_{j}$
$\hat{\mathbf{C}}$ The pseudo ground truth $\mathbf{v}^{m}_{r,j}$ A keyframe on voxel space get through $\Theta^{m}_{p}$
$N^{d}$ The number of skills in a continual skill learning task $\mathbf{Y}^{m}_{j}$ The ground truth for prediction the next keyframe
$N^{b}$ The number of voctors in $\mathcal{B}$ $\hat{\mathbf{Y}^{m}_{j}}$ The soft label extracted by the perceiver model $\Theta^{m-1}_{p}$
$N^{s}$ The number of learned skills $\mathbf{X}$ The input feature for projection layers in Transformer
$N^{l}$ The learnable latent vector quantity $D$ The dimension of latent vectors in projection layers in Transformer
$N^{p}$ The number of encoded embeddings of input data $D^{s}$ The dimension of $\mathbf{l}_{s}$
$N^{e}$ The number of encoded embeddings of the input language instruction
$N^{r}$ The size of LoRA layers
A.1 Dataset Description
In order to emulate the operational environments of domestic robots, we design two NBRL benchmark datasets using RLBench, called Kitchen and Living Room. In particular, we assemble Kitchen by gathering 10 skills that are pertinent to kitchen settings, and we construct Living Room, which includes 12 skills that are related to living room situations. Each skill includes a training set of 20 episodes and a test set of 25 episodes. We provide examples of brief overviews of these skills and the number of key frames for each skill in Tab. 6 and visualization in Figs. 4-5. In addition, the No. represents the corresponding skill order mentioned in the Experiment section. The Avg. Kf describes the average number of key frames for the skill over 20 episodes in training set.
Table 6: Skill Details in Kitchen & Living Room
No. Skill in Kitchen Language Example Avg. Kf
1 close microwave "close microwave" 2.3
2 meat off grill "take the steak off the grill" 6.0
3 open grill "open the grill" 4.5
4 open wine bottle "open wine bottle" 3.3
5 pick up cup "pick up the red cup" 3.8
6 turn tap "turn left tap" 3.0
7 place wine "stack the wine bottle to the left of the rack" 6.1
8 put knife on "put the knife on the chopping board" 5.1
9 stack wine "stack wine bottle" 7.4
10 take plate off "take plate off the red colored rack" 6.3
No. Skill in Living Room Language Example Avg. Kf
1 close door "close the door" 4.8
2 close laptop lid "close laptop lid" 6.0
3 hang frame "hang frame on hanger" 5.3
4 lamp on "turn on the light" 3.5
5 open drawer "open the bottom drawer" 4.8
6 open window "open left window" 6.0
7 push buttons "push the white button" 4.8
8 put item in drawer "put the item in the top drawer" 14.3
9 put rubbish in bin "put rubbish in bin" 5.0
10 sweep to dustpan "sweep dirt to the tall dustpan" 6.2
11 take usb out "take usb out of computer" 3.3
12 water plants "pour some water on the plant" 6.3
Figure 4: Visualization of manipulation skills in dataset Kitchen.
Figure 5: Visualization of manipulation skills in dataset Living Room.
Table 7: Hyper-parameters used in NBAgent
Variable Name Value
image size 128 × 128 × 3
input voxel size 100 × 100 × 100
batch size $B$ 2
optimizer LAMB
learning rate 0.0005
number of transformer blocks 6
number of sampled points for NBAgent 64
number of latents in Perceiver Transformer 2046
dimension of Stable Diffusion features 512
dimension of CLIP language features 512
hidden dimension of NeRF blocks 512
size of LoRA layers $N^{r}$ 10
base task iterations 80K
incremental task iterations 20K
SSR loss weight $\lambda_{1}$ 0.1
SSR loss function $\mathcal{L}_{SSR}$ MSE-loss
SRD loss weight $\lambda_{2}$ 0.2
SRD loss function $\mathcal{L}_{SRD}$ KL divergence
SRD temperature $\mathcal{T}$ 3
novel skill threshold $\delta$ 0.8
size of memory buffer $\mathcal{M}$ 4
number of base skills in living room 6
number of novel skills in living room 3, 2
number of base skills in kitchen 5
number of novel skills in kitchen 1, 5
Table 8: Comparison results on Kitchen under the setting of 5-5.
Comparison Methods 1 2 3 4 5 6 7 8 9 10 All. Imp.
PerAct 96.0 77.3 53.3 30.7 37.3 52.0 2.7 18.7 4.0 73.3 44.5 $\Uparrow$ 0.4
GNFactor 92.0 60.0 70.7 12.0 46.7 52.0 5.3 22.7 10.7 70.7 44.3 $\Uparrow$ 0.6
Fine-Tuning 10.7 0.0 33.3 33.3 1.3 49.3 6.7 14.7 5.3 56.0 21.1 $\Uparrow$ 23.8
ER 96.0 72.0 56.0 18.7 37.3 45.3 0.0 14.7 1.3 66.7 40.8 $\Uparrow$ 4.1
Ours-w/oSEP&SRD 86.7 74.7 64.0 28.0 26.7 52.0 6.7 17.3 8.0 49.3 41.3 $\Uparrow$ 3.6
Ours-w/oSRD 8.0 81.3 36.0 38.7 41.3 62.7 4.0 46.7 9.3 68.0 39.6 $\Uparrow$ 5.3
[…]
Table 9: Comparison results on Living Room under the setting of 6-2.
Comparison Methods 1 2 3 4 5 6 7 8 9 10 11 12 All. Imp.
PerAct 5.3 66.7 0.0 84.0 38.7 13.3 16.0 2.7 41.3 0.0 97.3 6.7 31.0 $\Uparrow$ 9.6
GNFactor 2.7 92.0 1.3 84.0 80.0 28.0 8.0 1.3 46.7 32.0 100 0.0 40.0 $\Uparrow$ 0.6
Fine-Tuning 0.0 37.3 0.0 0.0 0.0 0.0 0.0 0.0 0.0 0.0 96.0 18.7 12.7 $\Uparrow$ 27.9
ER 4.0 89.3 17.3 80.0 38.7 14.7 20.0 2.7 60.0 4.0 90.7 4.0 35.4 $\Uparrow$ 5.2
Ours-w/oSEP&SRD 2.7 82.7 14.7 76.0 30.7 24.0 18.7 9.3 76.0 4.0 94.7 10.7 37.0 $\Uparrow$ 3.6
Ours-w/oSRD 12.0 65.3 14.7 82.7 49.3 18.7 28.0 18.7 81.3 6.7 100.0 0.0 39.8 $\Uparrow$ 0.8
[…]
A.2 Implementation Details
As shown in Tab. 7, we give the hyper-parameters used in NBAgent. One agent is trained with two NVIDIA A40 GPU for 2 days (80K + 40K iterations under the 6-3 setting in Living Room) with a batch size of 2. The overall learning rate of the training process is 0.0005, optimized by the LAMB algorithm. All methods employ the same methodology to construct a voxel grid of size $100^{3}$. During the NeRF training, we utilize an additional 19 camera perspectives to furnish supervisory information. For NBRL, we conduct two NBRL settings in each of the two scenarios, which are 5-5, 5-1 in Kitchen, and 6-3, 6-2 in Living Room.
A.3 More Comparison Experiments
Tabs. 8 and 9 show the detailed scores of each model in two scenarios under the NBRL settings of 5-5 and 6-2.With the configuration of memory buffer size $|\mathcal{M}|=4$, the average score of NBAgent across all skills still surpasses other models, approximately between $0.4\%\sim 23.8\%$ and $0.6\%\sim 27.9\%$. This indicates that NBAgent performs the best in solving NBRL problems. Fig. 6 shows the visualization results of more skills among various models. We observe that some skills are not difficult to learn the general actions, but when the skill requires the target object to be in a specific location, the agent has difficulty completing it accurately; for example, ’open the drawer’ is not difficult, but when the requirement is ’open the top drawer’, the performance of NBAgent is better than other models.
Figure 6: Visualization of prediction results on various skills between Ours, Ours-w/oSRD, Ours-w/oSEP&SRD, ER and Fine-Tuing.
A.4 Limitation
While we have found that NBAgent achieves state-of-the-art results on the Never-ending Behavior-cloning Robot Learning (NBRL) problem, we have also identified the following limitations: 1. NBAgent collects keyframes as learning targets using predefined rules in the demonstration dataset, but these rules are not universally applicable to all manipulation tasks. 2.NBAgent cannot continuously learn skills that require continuous actions as it predicts actions based on time-discrete control.
A.5 Social Impact
In practical multi-task robot learning, we explore the Never-ending Behavior-cloning Robot Learning (NBRL) problem and propose a novel NBAgent. The goal is to continually learn new skills in an incremental manner, allowing for unique customized operational abilities. For example, robots can learn new skills by imitating specific demonstration datasets tailored to the owner’s needs. Additionally, we address catastrophic forgetting in continual skill learning by transferring skill-shared semantic of 3D scenes using a skill-shared semantic rendering module and a skill-shared representation distillation module.
Typically, robots need strong learning capabilities to adapt to various work scenarios effectively. Having the ability for continual learning allows robots to provide better user experiences. Furthermore, this approach enables robots to learn specific operational skills based on the diverse needs of different user groups. For instance, household robots need to learn operational tasks in various contexts. Therefore, the proposed Never-ending Behavior-cloning Robot Learning problem is worth investigating.

### v3

@@ IV-B Comparison Performance on RLbench
We present comparative results between NBAgent and baseline methods on Kitchen and Living Room datasets in Tabs. II-VI and Fig. 3. As shown in Tab. II, NBAgent significantly outperforms compared methods by $1.5\%\sim 41.4\%$ in terms of Avg. metric on two Kitchen dataset settings and $1.3\%\sim 19.3\%$ on two Living Room dataset settings. Additionally, NBAgent surpasses the SOTA method M2Distill [49] by $1.5\%\sim 2.4\%$ on the Kitchen settings and $1.8\%\sim 2.7\%$ on the Living Room settings in terms of the Avg. metric. These results indicate that NBAgent effectively addressing catastrophic forgetting of old skills through skill-specific and skill-shared knowledge learning. From the comparative analysis, we derive the following observations: 1) Superiority to Learn Novel Skills: Despite observing the skill data streams, NBAgent surprisingly surpasses the joint training methods PerAct [53] and GNFactor [70] by $0.4\%\sim 1.7\%$ in average success rate across all skills, providing additional evidence to support the efficacy of our model. 2) Robust Forgetting Mitigation: Throughout the whole NBRL process, our NBAgent exhibits $1.9\%\sim 6.7\%$ lower forgetting metrics than all comparison methods on Living Room 6-2 setting, highlighting its robustness in retaining learned skills.
In addition, we illustrate key simulation inference frames as qualitative results in Fig. 3. We can find that comparison methods frequently fail to complete manipulation tasks due to two critical limitations: 1) inability to accurately recognize and localize objects, and 2) suboptimal action cloning caused by catastrophic forgetting. In contrast, NBAgent addresses these challenges better, confirming its effectiveness in preserving skill-specific and skill-shared knowledge.
IV-C Skill-Wise Comparison Performance
We presented the skill-wise comparative results in Tabs. IV, V, III and VI. The experimental results reveal that our model achieves the highest average skill-wise success rate, demonstrating improvements of $2.1\%\sim 36.3\%$ on two Kitchen dataset settings and $2.2\%\sim 25.6\%$ on two Living Room dataset settings, compared to other methods. These findings further validate the effectiveness of our framework in addressing the challenges of the NBRL problem. NBAgent achieves superior performance over baseline methods in most skill-wise comparisons, highlighting its capability to mitigate catastrophic forgetting through dual mechanisms: 1) Skill-Shared Knowledge Preservation: The proposed SSR and SRD modules facilitate effective transfer of 3D scene semantics between between old and current models. 2) Skill-Specific Knowledge Adaptation: The SEP module enables continual acquisition of novel skills while maintaining skill-specific attribute.
@@ IV-C Skill-Wise Comparison Performance
TABLE IX: Comparison results in terms of success rate (%) on Living Room dataset when setting the various size of memory buff $\mathcal{M}$.
@@
Figure 8: Analysis of hyper-parameters $\lambda_{1}$ and $\lambda_{2}$ under the 5-5 setting.
@@
TABLE X: Comparison results in terms of success rate (%) on Living Room dataset when setting the various task orders.
Task order 6-3 $(2~\mathrm{steps})$ 6-2 $(4~\mathrm{steps})$
1-6 7-12 All Avg. For. 1-6 7-12 All Avg. For.
Order 1 44.9 42.2 43.6 47.6 7.7 45.8 35.3 40.6 43.5 10.9
Order 2 41.4 43.5 42.5 46.1 8.4 40.6 38.5 40.0 41.8 8.6
Order 3 45.6 42.1 43.9 48.4 7.5 45.9 37.9 41.9 44.0 8.5
IV-D Ablation Studies
Ablation on skill-shared attribute. Compared to Ours, the scores of Ours-w/oSRD on old skills are dropped by $3.8\%\sim 12.5\%$ (i.e., 1-5 and 1-6 in Tab. II). This indicates that SRD can effectively learn skill-shared knowledge to tackle semantic overlooking on old skills, To evaluate the effectiveness of SSR module, we visualize the rendering results in Fig. 7. It suggests that SSR module can efficiently transfer skill-shared semantics of 3D scenes, thereby achieving an improvement about $0.5\%\sim 3.1\%$ in terms of Avg compared Ours-w/oSEP&SRD with ER in Tab. II. The above mentioned demonstrates that NBAgent can significantly tackle old skill forgetting by preserving skill-shard semantics of 3D scenes and voxel representation.
Figure 9: Analysis of hyper-parameters $\delta$ under the 5-5 setting.
TABLE XI: Ablation studies on the external models in SSR module under the settings of 5-5, where the external models including CLIP [45], DINO [7] and SD (Stable Diffusion) [48].
Variants Base Base+DINO Base+CLIP   Ours
Modules DINO
CLIP
SD
5-5 $(2~\mathrm{steps})$ 1-5 49.4 50.4 52.8 53.6
6-10 34.8 34.6 36.0 36.3
ALL 42.1 42.5 44.4 44.9
Avg. 49.4 50.2 51.6 52.5
For. 10.6 10.8 8.2 6.4
TABLE XII: Comparison results in long-term learning under the setting of 5-1 (12 steps).
Comparison Methods 1-5 6-16 All Avg. For.
Fine-Tuning 1.6 9.1 6.8 9.2 58.4
iCaRL [46] (CVPR’2017) 36.0 28.6 30.9 36.4 26.5
PODNet [14] (ECCV’2020) 36.8 26.4 29.7 35.2 28.9
C-LoRA [54] (TMLR’2024) 41.6 30.4 33.9 38.7 21.6
M2Distill [49] (ICRA’2025) 40.8 36.8 38.1 41.7 18.5
Ours 43.6 40.4 41.4 44.5 15.3
TABLE XIII: Training parameters, memory cost, training time per 100 iterations and comparison results in terms of All. under the setting of 6-3.
Comparison Methods Trainable Memory Training All
Parameters Storage Time (100 it/s)
Fine-Tuning 48.3M 0.2G 51.2s 17.1
ER [8] (ECCV’2018) 48.3M 5.1G 56.4s 37.4
LWF [35] (TPAMI’2017) 48.3M 5.1G 78.6s 36.3
iCaRL [46] (CVPR’2017) 48.3M 5.1G 81.4s 37.0
PODNet [14] (ECCV’2020) 48.3M 5.1G 72.7s 38.1
C-LoRA [54] (TMLR’2024) 18.1M 4.8G 60.6s 40.7
M2Distill [49] (ICRA’2025) 48.3M 4.8G 92.3s 41.1
Ours 18.1M 4.8G 76.4s 43.6
Ablation on skill-specific attribute. To evaluate effectiveness of our NBAgent on learning skill-specific knowledge, we eliminate our proposed SEP module and present results in Tabs. II, IV and V. Ours-w/oSRD outperforms Ours-w/oSEP&SRD on novel skills by $1.3\%\sim 11.4\%$, (i.e., 6-10 and 7-12 in Tab. II), which demonstrates that SEP benefits our model in learning novel skills by performing skill-specific knowledge learning.
Ablation on memory size. We also explore the impact of various sizes of memory buffer $\mathcal{M}$. As shown in Tab. IX, with $|\mathcal{M}|=6$, the forgetting rate of our model is notably reduced by $1.1\%\sim 10.8\%$ compared to $|\mathcal{M}|=4$ and $2$, respectively. These results indicate that increasing the memory size substantially mitigates catastrophic forgetting, while introducing increased memory overhead.
Ablation on the external models in SSR module. To assess the effectiveness of the SSR module’s reliance on external models, we conducted ablation studies with Base (NeRF-only rendering), Base+DINO [7], and Base+CLIP [45]. As shown in Tab. XI, using only NeRF captures the geometric structure but fails to preserve the skill-shared semantics of the 3D scene, resulting in degraded task performance. Replacing Stable Diffusion with DINO or CLIP transfers only partial semantic knowledge, yet still falls short of the full SSR module. These findings underscore the crucial role of Stable Diffusion in ensuring robust 3D skill-shared semantic consistency.
Ablation on hyper-parameters in NBAgent. We analyze the influence of the trade-off parameters $\lambda_{1}$ and $\lambda_{2}$ on the overall performance under the 5-5 setting. As illustrated in Fig. 8, both the “All” and “Avg.” metrics exhibit a similar trend, where the optimal performance is achieved when $\lambda_{1}$ = 0.1 and $\lambda_{2}$ = 0.2. In addition, we compute the average similarity $\mathcal{C}_{max}$ using the data from incremental tasks under the 5+5 setting and present the intra-skill and inter-skill similarities in Fig. 9. The results indicate that the hyperparameter $\delta=0.8$ effectively distinguishes between old skills (with higher similarity) and new skills (with lower similarity).
Ablation on task order. To evaluate the effect of task order on continual learning performance, we conduct additional experiments using three different randomized task orders. The results in Tab. X show that NBAgent maintains stable performance across different task orders, indicating its robustness to task sequence variations. The performance fluctuations among orders are minor (within about 2%), suggesting that catastrophic forgetting is effectively mitigated, which demonstrates that NBAgent achieves consistent NBRL regardless of the training order.
IV-E Analysis of Task Complexity
As shown in Tab. XII, we present the comparison results between NBAgent and other methods. NBAgent consistently outperforms the baselines by $2.8\%\sim 35.3\%$ in terms of the average metric. These results demonstrate that NBAgent effectively mitigates catastrophic forgetting of previously learned skills through the joint learning of skill-specific and skill-shared knowledge in long-term learning. Additionally, as shown in Fig. 6, we report the forgetting rates of NBAgent and the compared baselines. The forgetting rate of our approach increases more gradually than that of other methods, demonstrating its stronger resistance to catastrophic forgetting and its ability to retain learned knowledge more stably over time.
IV-F Analysis of Computation Complexity
In this subsection, we analyze the computational complexity of our framework in addressing the NBRL problem. As summarized in Tab. XIII, we present comparative analyses with baseline methods across two critical metrics: trainable parameters and memory storage. To ensure equitable evaluation conditions, all experiments are conducted on identical hardware settings utilizing two NVIDIA A6000 GPUs. From Tab. XIII, our framework achieves state-of-the-art parameter efficiency through the proposed SEP module. By only training the skill-specific latent space in Eq. (12) and skill-specific LoRA layers in Eq. (13), the model reduces $62.6\%$ trainable parameters of the total parameters and achieve the competitive efficiency in training time per 100 iterations, while reaching the highest success rates. These results demonstrate that our model achieves an excellent optimal trade-off between computational efficiency and task effectiveness.
V Conclusion and Limitation
@@
Conclusion. In this paper, we systematically investigate a pioneering Never-ending Behavior-cloning Robot Learning (NBRL) problem and propose NBAgent, a continual learning framework to continually learn skill-wise knowledge. Specifically, we design a skill-shared semantic rendering module and a skill-shared representation distillation module to tackle 3D scene representation overlooking on past encountered skills. Meanwhile, we propose a skill-specific evolving planner to decouple the skill-wise knowledge through adaptive latent space partitioning and parameter-efficient LoRA layers, enabling simultaneous effective learning of novel skills and mitigation of catastrophic forgetting. We develop two NBRL benchmarks to demonstrate the manipulation accuracy, which verifies the effectiveness of our NBAgent against baselines. Through extensive experiments on robotic manipulation tasks, we validate that NBAgent achieves state-of-the-art performance in NBRL problem.
Limitation. Although NBAgent achieves state-of-the-art results on the Never-ending Behavior-cloning Robot Learning (NBRL) problem, the primary limitation persists: The current implementation relies on predefined rules for keyframe extraction from demonstration datasets. While effective for constrained manipulation tasks (e.g., drawer opening, object grasping), these heuristics may not generalize to scenarios requiring dynamic skill segmentation, such as fluid pouring or deformable object manipulation.